# JWST Insights into Narrow-line Little Red Dots

### v1

@@ 4.1.2 Comparison of UV-based and H-based SFRs
We report the and values in Table 2. The comparisons of the UV-based and H-based SFRs with and without dust correction for LRDs are shown in Figure 11, along with those of the star-forming galaxies from K. Nakajima et al. (2023) and M. Curti et al. (2024) measured by the same method. The and are consistent for star-forming galaxies in both the dust-uncorrected and dust-corrected cases, indicating modest and similar levels of dust extinction affecting their UV and H emission. In contrast, for LRDs, including narrow-line LRDs, is significantly higher than when no dust correction is applied, but the two become consistent after the dust correction. This result suggests that UV continuum and narrow H emission in LRDs are likely dominated by star formation, which supports the purely star-forming galaxy scenario, whereas the broad component may originate from a different mechanism (see discussion in M. Kokubo & Y. Harikane, 2024; R. Maiolino et al., 2025, for possible origins).
@@
and can be used to calculate the burstiness parameter that is an indicator of the current burst state of a galaxy (A. Broussard et al., 2019). The average parameter of the narrow-line LRDs is nearly 0, implying that the galaxy population does not have a rapidly rising or falling star formation recently. The scatters of the parameter are similar for narrow-line LRDs and normal star-forming galaxies, indicating their similar bursty levels.
@@
In this purely star-forming galaxies scenario, the inferred SFRs are quite high (up to , see Figure 11) for broad-line LRDs, which is contradicted to the lack of strong IR and [C ii] emission for most LRDs, as revealed by deep ALMA/NOEMA observations (I. Labbé et al., 2023; C. C. Williams et al., 2024; K. Chen et al., 2025; C. M. Casey et al., 2025; M. Xiao et al., 2025; D. J. Setton et al., 2025). These observations have primarily targeted relatively bright LRDs with broad-line detections, so it remains unclear whether the same trend applies to our narrow-line LRDs. Moreover, narrow-line LRDs tend to have more moderate SFRs (typically ), and their expected IR emission is correspondingly weaker.
4.1.3 Balmer Decrement
@@
In our sample, there are eight LRDs (including two narrow-line LRD UID-8219 and UID-12329) with H line detections in the high- or medium-resolution spectra with . In the galaxy-only scenario, we further check the consistency of the stellar extinction derived from the SED fitting and the nebular extinction from the line flux ratio of H and H, i.e., the Balmer decrement. For broad-line LRDs, we only use their narrow-line components. The detailed calculation can be found in Appendix A. The comparison of the two measurements is shown in Figure 12. For the two narrow-line LRDs, their and values are roughly consistent, suggesting that the galaxy-only SED fitting results are reasonable. In contract, broad-line LRDs tend to have larger , indicating that the galaxy-only SED model is not sufficient to describe these sources. The values of two LRDs are even smaller than zero. This is because they have H/H line ratios smaller than 2.86, which is different from the Case B assumption. Similar results have been reported by previous works (e.g., N. Pirzkal et al., 2024; B. Sun & H. Yan, 2025).
4.2 AGNs with Low Black Hole Masses
@@
A large fraction of LRDs are believed to be AGNs, as indicated by the presence of broad Balmer lines. The observed V-shape continuum may result from AGN emission attenuated by very dense, dust-free gas, possibly combined with a scattered or leaked AGN component or host galaxy light (e.g., J. E. Greene et al., 2024; K. Inayoshi & R. Maiolino, 2025; X. Ji et al., 2025; R. P. Naidu et al., 2025). Alternatively, the V-shape continuum may be due to a UV-flattened dust extinction law (e.g., M. Killi et al., 2024; Z. Li et al., 2025). Recently, it is proposed that the continuum shape may be intrinsic, perhaps reflecting a binary massive black hole system or a two-component accretion disk structure (e.g., K. Inayoshi et al., 2025; C. Zhang et al., 2025). If we believe that V-shape continuum is a unique feature of a special population of AGN, narrow-line LRDs may represent a population of AGNs at the very low-mass end of the SMBH mass distribution, which could naturally lead to the absence of broad Balmer lines. The excesses in and of the narrow-line LRDs compared to typical star-forming galaxies (Figure 6) also suggest an additional contribution to the H emission beyond that from star formation, possibly originating from a central AGN. We therefore consider a scenario in which these narrow-line LRDs with V-shape continua are AGNs hosting low-mass SMBHs, leading to intrinsically narrow Balmer lines, which is similar to the case of narrow-line Seyfert 1 galaxies (NLS1s; e.g., D. Grupe & S. Mathur, 2004; J. K. Williams et al., 2018). This scenario is also in line with the recently proposed interpretation that the observed widths of the broad lines in high-redshift JWST AGN and LRDs are due to electron or Balmer scattering of intrinsic narrow lines by medium located outside the broad line regions rather than due to the motions of BLR clouds (R. P. Naidu et al. 2025; V. Rusakov et al. 2025, but see I. Juodžbalis et al. 2025 for a debate). Narrow-line LRDs may represent a special population in which scattering is absent, allowing the intrinsic narrow line to be directly observed.
4.2.1 Black Hole Mass Measurement
@@
We estimate the black hole masses of the narrow-line LRDs using the local virial relations, which link the black hole mass with the continuum/line luminosities and line widths. We use the relation calibrated by A. E. Reines et al. (2013):
with a scaling factor following A. E. Reines & M. Volonteri (2015). We adopt and measured from the single narrow Gaussian fitting. The H luminosity is dust-corrected as , where the ranges from 1.7 to 3.1. The values here are derived from the best-fit pure galaxy SED model described in Section 4.1, but they are consistent with typical estimates for broad-line LRDs reported in other studies (e.g., J. E. Greene et al., 2024; D. D. Kocevski et al., 2024). The impact of the adopted luminosity on the final black hole mass is significantly smaller than that of the line width. We use the observed , which can be regarded as the upper limit of the intrinsic . In this way, we attribute the entire H line to the AGN contribution. In reality, the host galaxy also contributes to the H line. In addition, the actual black hole mass would be even smaller if the sources are dust-free and no significant extinction correction is needed (e.g., R. P. Naidu et al., 2025; K. Inayoshi et al., 2025). Therefore, the derived black hole masses are conservative upper limits. The resultant black hole mass upper limits are listed in Table 2, with values in the range of . These values are much smaller than that of the typical JWST AGNs and broad-line LRDs (e.g., Y. Harikane et al., 2023; R. Maiolino et al., 2024), mainly due to the narrow H line widths of our sources.
4.2.2 – Relation
@@
We study these narrow-line LRDs in the context of the – relation. Due to the limited understanding of the unresolved central source, the results are open to multiple interpretations. The first possibility is that the emission from the unresolved point source mainly arises from extremely compact galaxies, as discovered by some works (e.g., A. C. Carnall et al., 2023; Y. Ono et al., 2023). This is also consistent with the lack of significant AGN signatures in the narrow-line LRDs. We adopted the largest possible stellar masses estimated in Section 4.1. The corresponding black hole mass versus stellar mass relation of the arrow-line LRDs is shown as the red stars in Figure 13. For comparison, we also plot the local () scaling relation (J. E. Greene et al., 2020; M.-Y. Zhuang & L. C. Ho, 2023) as well as other high-redshift Type 1 AGNs discovered with JWST (Y. Harikane et al., 2023; R. Maiolino et al., 2024; C.-H. Chen et al., 2025b). We find that three of the narrow-line LRDs are roughly consistent with or slightly below the local relation. The other two narrow-line LRDs likely host overmassive black holes, but their are still smaller than those in other JWST Type 1 AGNs.
@@
Another possibility is that the central point sources are dominated by AGN, while the possible extended emission contributed by the host galaxies is weak. In this case, we adopt the morphological fitting result using a central point source and a Sérsic component in Section 3.2. Three narrow-line LRDs can be well fitted by a single PSF model. The other two (UID-8219 and UID-12329) exhibit obviously extended emission. For UID-12329, the extended emission is significantly off the center. Although such off-centered extended emission in LRDs may originate from purely nebular emission photoionized by the LRD itself (C.-H. Chen et al., 2025a), we assume that it is from its host galaxy here.
@@
We perform SED modeling using CIGALE with the same configuration in Section 4.1.1 for the extended emission of UID-8219 and UID-12329. We add a relative error of 15% in quadrature to the uncertainties of the fluxes to account for the uncertainties introduced by the image fitting. The best-fit SED result of UID-12329 is shown in Figure 9(b), with the SED of the PSF component shown for comparison. The derived stellar masses are substantially lower than those inferred assuming the entire emission arises from the host galaxy. The red rest-optical continuum is dominated by a central point source, whereas the extended emission appears significantly bluer. This color contrast is consistent with the results of C.-H. Chen et al. (2025b) for broad-line LRDs and suggests minimal dust extinction for the host galaxies, which in turn leads to lower stellar mass estimates. We also use the observed to calculate the black hole mass of UID-8219 and UID-12329. The results are shown as the yellow stars in Figure 13. In this case, the black hole masses are not as overmassive as those in other JWST AGNs but still lie above the local relations. If dust corrections are needed for , these objects would shift further away from the local scaling relations.
4.3 X-ray Non-detection
@@
Previous X-ray stacking analyses of broad-line LRDs show that they have weak X-ray emission (e.g., T. T. Ananna et al., 2024; M. Yue et al., 2024; R. Maiolino et al., 2025), which may be attributed to radiative signatures of super-Eddington accreting black holes (e.g., P. Madau & F. Haardt, 2024; K. Inayoshi et al., 2024). We check the archival X-ray observations of the five narrow-line LRDs and find that they are all covered by the AEGIS-X Deep survey (K. Nandra et al., 2015) and the X-UDS survey (D. D. Kocevski et al., 2018), respectively. However, none of them has an individual X-ray detection. We use CSTACK333https://lambic.astrosen.unam.mx/cstack/. and follow the method in M. Yue et al. (2024) to stack the X-ray observations and compare the results with the expected count rate using the relation in C. Jin et al. (2012), assuming the column densities . The stacking still yields a non-detection with a total exposure time of 2.2 Ms. The comparison of the observed upper limits of the count rates between the expected values from the observed H luminosities for the 0.5–2 keV and 2–8 keV band is shown in Figure 14. The soft band provides tighter constraints than the hard band. However, in both cases, the constraints for both the individual sources and the stacked sample are not stringent enough. The current depth of the X-ray data does not allow us to determine whether narrow-line LRDs are X-ray weak.
5 Discussion
5.1 Comparison with Other Works
@@
We compare our LRD selection result with other studies, specifically V. Kokorev et al. (2024a) and D. D. Kocevski et al. (2024). Cross-matching the compact sources in our parent sample with the V. Kokorev et al. (2024a) LRD sample yields 38 common sources, of which 29 meet our LRD criteria. The remaining 9/38 () sources all exhibit very strong H, [O iii], and H emission lines, which boost their optical colors to be very red. After correcting for the emission line contribution (Section 3.1), their rest-frame optical continuum slopes fall outside the LRD criteria. Similarly, cross-matching with D. D. Kocevski et al. (2024) yields 61 common sources, with 50 classified as LRDs by our selection. Most of the remaining 11/61 () sources are also misclassified due to the strong emission lines, and two sources are due to the wrong photo-.
5.2 Velocity Dispersion and Host Dynamical Mass

### v2

@@ IV.1.2 Comparison of UV-based and H-based SFRs
We report the and values in Table 2. In our analysis here, the dust attenuation is derived from the CIGALE SED fitting in Section IV.1.1. For each source, we use the best-fit color excess for nebular lines and convert it to . Dust-corrected H luminosities are then calculated using . The UV dust correction is computed by the model SEDs provided by CIGALE: we evaluate the model flux density at the rest-frame NUV pivot wavelengths before and after the attenuation correction to obtain dust-corrected UV luminosities. The comparisons of the UV-based and H-based SFRs with and without dust correction for LRDs are shown in Figure 11, along with those of the star-forming galaxies from 2023ApJS..269...33N and 2024A&A...684A..75C measured by the same method. The and are consistent for star-forming galaxies in both the dust-uncorrected and dust-corrected cases, indicating modest and similar levels of dust extinction affecting their UV and H emission. In contrast, for LRDs, including narrow-line LRDs, is significantly higher than when no dust correction is applied, but the two become consistent after the dust correction. This result suggests that UV continuum and narrow H emission in LRDs are likely dominated by star formation, which supports the purely star-forming galaxy scenario, whereas the broad component may originate from a non-AGN mechanism, such as Raman scattering (see discussion in 2024arXiv240704777K; 2025MNRAS.538.1921M, for possible origins).
@@
and can be used to calculate the burstiness parameter that is an indicator of the current burst state of a galaxy (2019ApJ...873...74B). The average parameter of the narrow-line LRDs is nearly 0 if use the dust-corrected results, implying that the galaxy population does not have a rapidly rising or falling star formation recently. The scatters of the parameter are similar for narrow-line LRDs and normal star-forming galaxies, indicating their similar bursty levels.
@@
In this purely star-forming galaxies scenario, the inferred SFRs are quite high (up to , see Figure 11) for broad-line LRDs, which is contradicted to the lack of strong IR and [C ii] emission for most LRDs, as revealed by deep ALMA/NOEMA observations (2023Natur.616..266L; 2024ApJ...968...34W; 2025arXiv250522600C; 2025arXiv250518873C; 2025arXiv250301945X; 2025arXiv250302059S). These observations have primarily targeted relatively bright LRDs with broad-line detections, so it remains unclear whether the same trend applies to our narrow-line LRDs. Moreover, narrow-line LRDs tend to have more moderate SFRs (typically ), and their expected IR emission is correspondingly weaker.
IV.1.3 Balmer Decrement
@@
In our sample, there are eight LRDs (including two narrow-line LRD UID-8219 and UID-12329) with H line detections in the high- or medium-resolution spectra with . In the galaxy-only scenario, we further check the consistency of the stellar extinction derived from the SED fitting and the nebular extinction from the line flux ratio of H and H, i.e., the Balmer decrement. For broad-line LRDs, we only use their narrow-line components. The detailed calculation can be found in Appendix A. The comparison of the two measurements is shown in Figure 12. For the two narrow-line LRDs, their and values are roughly consistent, suggesting that the galaxy-only SED fitting results are reasonable. In contract, broad-line LRDs tend to have larger , indicating that the galaxy-only SED model is not sufficient to describe these sources. The values of two LRDs are even smaller than zero. This is because they have H/H line ratios smaller than 2.86, which is different from the Case B assumption. Similar results have been reported by previous works (e.g., 2024ApJ...969...90P; 2025arXiv250321896S).
IV.2 AGNs with Low Black Hole Masses
@@
A large fraction of LRDs are believed to be AGNs, as indicated by the presence of broad Balmer lines. The observed V-shape continuum may result from AGN emission attenuated by very dense, dust-free gas, possibly combined with a scattered or leaked AGN component or host galaxy light (e.g., 2024ApJ...964...39G; 2025ApJ...980L..27I; 2025arXiv250113082J; Lin2025_locallrd; LiuH_2025; 2025arXiv250316596N; Zhang2025_lensedlrd). Alternatively, the V-shape may result from a UV-flattened dust-extinction law (e.g., 2024A&A...691A..52K; 2025ApJ...980...36L). It is also proposed that the continuum shape is intrinsic, perhaps reflecting a binary massive black hole system or a two-component accretion-disk structure (e.g., 2025arXiv250505322I; 2025arXiv251109278W; 2025arXiv250512719Z). If we believe that V-shape continuum is a unique feature of a special population of AGN, narrow-line LRDs may represent a population of AGNs at the very low-mass end of the SMBH mass distribution, which could naturally lead to the absence of broad Balmer lines. The excesses in and of the narrow-line LRDs compared to typical star-forming galaxies (Figure 6) also suggest an additional contribution to the H emission beyond that from star formation, possibly originating from a central AGN.
We therefore consider a scenario in which these narrow-line LRDs with V-shape continua are AGNs hosting low-mass SMBHs, leading to intrinsically narrow Balmer lines. Within this framework, two possibilities may account for the non-detection of a classical broad component: (1) the broad line is sufficiently broad but intrinsically weak and thus buried below the current noise level; or (2) the broad-line width is intrinsically so narrow that it is blended into the observed line profile, analogous to narrow-line Seyfert 1 galaxies (NLS1s; e.g., 2004ApJ...606L..41G; 2018MNRAS.480...96W). In the second scenario, the observed emission lines can arise from multiple components: the Balmer emission is partly contributed by an intrinsically narrow broad-line region (BLR) associated with a low-mass SMBH, while a conventional NLR is still present and produces narrower Balmer lines and forbidden lines such as [O iii]. As a result, the observed H emission can be enhanced relative to purely star-forming galaxies, which is consistent with the elevated H luminosities and line widths observed in narrow-line LRDs compared to typical star-forming systems (Figure 6). The second scenario is also in line with the recently proposed interpretation that the observed widths of the broad lines in high-redshift JWST AGN and LRDs are due to electron or Balmer scattering of intrinsic narrow lines by medium located outside the broad line regions rather than due to the motions of BLR clouds (2025arXiv250316596N; 2025arXiv250316595R, but see 2025arXiv250403551J for a debate). Narrow-line LRDs may represent a special population in which scattering is absent, allowing the intrinsic narrow line to be directly observed.
IV.2.1 Black Hole Mass Measurement
Motivated by the two possibilities discussed above, we estimate conservative upper limits on the black hole masses of the narrow-line LRDs by considering both possibilities. First, since the majority of LRDs exhibit broad emission lines, it remains plausible that these sources also host a weak broad component hidden below the current noise level or within the observed narrow emission line. Indeed, joint analyses of PRISM and G395M data report broad-line evidence in three of our targets, where the detection are mainly driven by the PRISM data owing to its higher sensitivity (Hviding2025). However, the uncertainties in such joint analyses remain large because of the low spectral resolution of PRISM and the substantial difference in wavelength-dependent resolution between the grating and prism. If such broad components exist, they would imply very small BH masses. To evaluate the maximal strength of a potential broad component that could remain undetected at the present G395M spectra, we simulate the broad line detection probability of AGNs with different and under the current G395M noise level and observed narrow emission line.
@@ IV.2.1 Black Hole Mass Measurement
We estimate the black hole masses of the narrow-line LRDs using the local virial relations, which link the black hole mass with the continuum/line luminosities and line widths. We use the relation calibrated by 2013ApJ...775..116R:
with a scaling factor following 2015ApJ...813...82R. We use this single-epoch BH mass calibration and the bolometric correction (Chen2025) to link , , , and . Once and are specified, the corresponding broad H component is fully determined. We use the best-fit single-Gaussian model of the observed H line as the narrow component. We then combine the broad and narrow components, inject random noise sampled from the rms of the observed spectrum, and perform the two kinds of Gaussian fitting described in Section III.2. For each – pair, we estimate the broad-line detection probability using 50 simulations. A broad line is considered to be detected if the model with both narrow and broad components has a lower BIC than the model with only the narrow component.
The simulated broad-line detection probabilities of the five narrow-line LRDs are presented in Figure 13. We also show the position of JWST AGNs and low-redshift quasars for comparison. The simulated result shows that the broad H lines of these sources should already have been detected if their and are similar to other JWST AGNs. At high Eddington ratio () region, the FWHM of the expected broad line could be , making them blended with the narrow line in some cases, similar to NLS1 in the local universe. In contrast, if is too low, the broad line could be very broad but too weak to be detected. The shaded region enclosed by the cyan dashed line (corresponding to large ) marks the parameter space where a broad H component would be clearly distinguishable from the observed narrow emission line and can therefore be ruled out. In this regime, the simulated broad component becomes sufficiently broader or stronger than the narrow component. Operationally, we flag a model as distinguishable when either (1) the expected broad-line FWHM exceeds the narrow-line FWHM by more than while remaining below , and its flux is at least of the narrow component; or (2) the broad-line FWHM is comparable to the narrow component (within ), but its flux is larger than the narrow-line flux. Therefore, the non-detection of broad line suggest that these narrow-line LRDs should have smaller and/or than other LRDs.
Assuming the Eddington ratio to be unity, the corresponding black hole mass of the ruled out region or the high broad-line detection probablity region can be regarded as an upper limit. The resultant black hole mass upper limits are in the range of –.
For the second scenario, the entire observed H line may arise from AGN activity if the BLR itself is intrinsically narrow. In this case, the observed line width can be treated as an upper limit on the intrinsic BLR velocity dispersion, while any contribution from the NLR or host galaxy would further reduce the true black hole mass. Under this assumption, we apply the same virial relation using the H luminosity and FWHM measured from the single-Gaussian fit. We therefore the use the intrinsic and the total H luminosity to estimate the black hole upper limit. This procedure yields black hole mass upper limits of –. These limits are conservative, as any host-galaxy contribution to H would further lower the inferred black hole masses.
Because both scenarios are physically plausible and cannot be distinguished with the current data, we adopt the more conservative constraint by taking, for each source, the larger of the two upper limits derived from the two scenarios. The resulting black hole mass upper limits are all , significantly smaller than those of typical JWST AGNs and broad-line LRDs. We do not account for dust extinction in our simulations, so the inferred upper limits on black hole mass could be underestimated if significant extinction is present. Nevertheless, the impact of the luminosity correction on is modest compared to that of the line width. In addition, LRDs are expected to exhibit only modest dust extinction (2025arXiv250518873C; 2025arXiv250522600C; 2025arXiv250302059S). Even adopting a maximal value of (e.g., 2025arXiv250522600C) would increase the black hole mass upper limits by only 0.3 dex.
IV.2.2 – Relation
@@
We study these narrow-line LRDs in the context of the – relation. Due to the limited understanding of the unresolved central source, the results are open to multiple interpretations. The first possibility is that the emission from the unresolved point source mainly arises from extremely compact galaxies, as discovered by some works (e.g., 2023Natur.619..716C; 2023ApJ...951...72O). This is also consistent with the lack of significant AGN signatures in the narrow-line LRDs. We adopted the largest possible stellar masses estimated in Section IV.1. The corresponding black hole mass versus stellar mass relation of the arrow-line LRDs is shown as the red stars in Figure 14. For comparison, we also plot the local () scaling relation (Greene2020; 2023NatAs...7.1376Z) as well as other high-redshift Type 1 AGNs discovered with JWST (2023ApJ...959...39H; 2024A&A...691A.145M; 2025ApJ...983...60C). We find that four of the narrow-line LRDs are roughly consistent with or slightly below the local relation. The remaining narrow-line LRD likely hosts an overmassive black hole, but its is still smaller than those in other JWST Type 1 AGNs.
@@
Another possibility is that the central point sources are dominated by AGN, while the possible extended emission contributed by the host galaxies is weak. In this case, we adopt the morphological fitting result using a central point source and a Sérsic component in Section III.2. Three narrow-line LRDs can be well fitted by a single PSF model. The other two (UID-8219 and UID-12329) exhibit obviously extended emission. For UID-12329, the extended emission is significantly off the center. Although such off-centered extended emission in LRDs may originate from purely nebular emission photoionized by the LRD itself (2025arXiv250503183C), we assume that it is from its host galaxy here.
@@
We perform SED modeling using CIGALE with the same configuration in Section IV.1.1 for the extended emission of UID-8219 and UID-12329. We add a relative error of 15% in quadrature to the uncertainties of the fluxes to account for the uncertainties introduced by the image fitting. The best-fit SED result of UID-12329 is shown in Figure 9(b), with the SED of the PSF component shown for comparison. The derived stellar masses are substantially lower than those inferred assuming the entire emission arises from the host galaxy. The red rest-optical continuum is dominated by a central point source, whereas the extended emission appears significantly bluer. This color contrast is consistent with the results of 2025ApJ...983...60C for broad-line LRDs and suggests minimal dust extinction for the host galaxies, which in turn leads to lower stellar mass estimates. The results are shown as the yellow stars in Figure 14. In this case, the black hole masses are not as overmassive as those in other JWST AGNs but still lie above the local relations. If dust corrections are needed for , these objects would shift further away from the local scaling relations.
IV.3 X-ray Non-detection
@@
Previous X-ray stacking analyses of broad-line LRDs show that they have weak X-ray emission (e.g., 2024ApJ...969L..18A; 2024ApJ...974L..26Y; 2025MNRAS.538.1921M), which may be attributed to radiative signatures of super-Eddington accreting black holes (e.g., 2024arXiv241000417M; 2024arXiv241203653I). We check the archival X-ray observations of the five narrow-line LRDs and find that they are all covered by the AEGIS-X Deep survey (2015ApJS..220...10N) and the X-UDS survey (2018ApJS..236...48K), respectively. However, none of them has an individual X-ray detection. We use CSTACK333https://lambic.astrosen.unam.mx/cstack/. and follow the method in 2024ApJ...974L..26Y to stack the X-ray observations and compare the results with the expected count rate using the relation in 2012MNRAS.420.1825J, assuming the column densities . The stacking still yields a non-detection with a total exposure time of 2.2 Ms. The comparison of the observed upper limits of the count rates between the expected values from the observed H luminosities for the 0.5–2 keV and 2–8 keV band is shown in Figure 15. The soft band provides tighter constraints than the hard band. However, in both cases, the constraints for both the individual sources and the stacked sample are not stringent enough. The current depth of the X-ray data does not allow us to determine whether narrow-line LRDs are X-ray weak.
V Discussion
V.1 Comparison with Other Works
@@
We compare our LRD selection result with other studies, specifically 2024ApJ...968...38K and 2024arXiv240403576K. Cross-matching the compact sources in our parent sample with the 2024ApJ...968...38K LRD sample yields 38 common sources, of which 29 meet our LRD criteria. The remaining 9/38 () sources all exhibit very strong H, [O iii], and H emission lines, which boost their optical colors to be very red. After correcting for the emission line contribution (Section III.1), their rest-frame optical continuum slopes fall outside the LRD criteria. Similarly, cross-matching with 2024arXiv240403576K yields 61 common sources, with 50 classified as LRDs by our selection. Most of the remaining 11/61 () sources are also misclassified due to the strong emission lines, and two sources are due to the wrong photo-.
V.2 Velocity Dispersion and Host Dynamical Mass